### v1

# Learning Visual Hierarchies with Hyperbolic Embeddings

Abstract
Structuring latent representations in a hierarchical manner enables models to learn patterns at multiple levels of abstraction. However, most prevalent image understanding models focus on visual similarity, and learning visual hierarchies is relatively unexplored. In this work, for the first time, we introduce a learning paradigm that can encode user-defined multi-level visual hierarchies in hyperbolic space without requiring explicit hierarchical labels. As a concrete example, first, we define a part-based image hierarchy using object-level annotations within and across images. Then, we introduce an approach to enforce the hierarchy using contrastive loss with pairwise entailment metrics. Finally, we discuss new evaluation metrics to effectively measure hierarchical image retrieval. Encoding these complex relationships ensures that the learned representations capture semantic and structural information that transcends mere visual similarity. Experiments in part-based image retrieval show significant improvements in hierarchical retrieval tasks, demonstrating the capability of our model in capturing visual hierarchies.
1 Introduction
Humans organize knowledge of the world into hierarchies [29] for efficient knowledge management. Developing models that encode such hierarchies is crucial for creating systems with holistic world understanding aligned with human perception. While this topic spans various modalities, we focus on encoding hierarchies in the visual domain. For many large image datasets, objects can be organized according to latent hierarchies [29], as evidenced by power law distributions [39]. However, most prevalent image understanding models [41, 17, 16, 5, 15] focus on preserving visual similarity, and consequently, learning hierarchies in the visual domain remains relatively unexplored. These limitations hinder models’ ability to generalize to tasks requiring hierarchical reasoning such as understanding scenes, object parts, and their interactions at multiple levels of abstraction. We illustrate in Fig. 1 an example that shows the complexity of a visual hierarchy where elements share similarities both within and across categories.
Recent works have demonstrated the utility of hyperbolic representation spaces for capturing hierarchical relationships in an unsupervised setting [8, 37, 35]. This emergent latent structure is appealing; nevertheless, meaningful hierarchies are often task and data dependent, and aligning model behavior with such human-defined hierarchies is essential for many applications. To this end, we introduce a learning paradigm that can encode user-defined multi-level hierarchies using hyperbolic space. Specifically, we first define a general part-based image hierarchy using object and part level annotations within and across images. Then, we introduce a model capable of structuring the latent space to preserve the defined hierarchy using only image/object level information. To our knowledge, we are the first to encode user-defined visual hierarchies without relying on explicit hierarchical labels or additional modalities. Finally, we introduce an evaluation metric to effectively measure hierarchical image retrieval.
Note that, hierarchy is an asymmetric relationship and has a high branching factor (see Fig. 1). To this end, we adopt the hyperbolic geometry as it provides a continuous approximation of such tree-like structures [29, 40]. To enforce the hierarchy, we break it into pairwise entailment relationships between images, objects, and parts, at multiple levels within an image as well as across images at category level. For pairwise entailment (asymmetric), we adapt the recently proposed angle-based asymmetric distance metric [37] within the contrastive learning paradigm, and extend it to handle cases with multiple positive relationships. In contrast to symmetric distances such as the inner-product used in [30, 21, 13, 33], this angle-based distance offers an additional degree of freedom along the radial axis to form emergent structures in the latent space.
Figure 1: An illustration of part-based image hierarchy organized in hyperbolic space. At the highest level, we see the urban environment, composed of buildings, streets, and sky. Zooming in, we find the building category, which further divides into skyscrapers, mid-rise structures, and more. Each of them has its own visual elements, which in turn can be decomposed into sub-elements. Best viewed when zoomed in.
For experimentation, we build a dataset of visual hierarchies using the bounding box annotations of OpenImages [25]. Our dataset includes entailment relationships between scenes, objects, and parts, within a single image as well as across images at the category level. Similarly, for hierarchical retrieval evaluation, we use the full training set to create ground truth hierarchy trees per scene/object. Additionally, we design a metric for evaluating hierarchical retrieval based on the optimal transport distance between the label distribution of the retrieval set and ground truth label distribution within the hierarchy tree. This, along with precision and recall metrics, demonstrates that our method captures semantic and structural information transcending mere visual similarity. Furthermore, we show the superior generalization capabilities of our model via zero-shot image classification experiments.
Our contributions can be summarized as follows:
  • To our knowledge, for the first time, we introduce a new learning paradigm to effectively encode user-defined multi-level visual hierarchies in hyperbolic space that does not require explicit hierarchical labels.
  • We adapt a contrastive loss using hyperbolic angle-based distance metric to enforce pairwise entailment relationships, and empirically demonstrate that pairwise entailment is sufficient to learn complex visual hierarchies.
  • We introduce an optimal transport based evaluation metric to measure hierarchical image retrieval performance.
  • We demonstrate superior generalization capabilities of our model beyond the user-defined hierarchies via zero-shot image classification evaluation and ablation experiments.
2 Preliminaries
We briefly review essential concepts in hyperbolic geometry here. We refer the reader to  [38] for a comprehensive treatment. Hyperbolic spaces are Riemannian manifolds with constant negative curvature and are fundamentally different from Euclidean or spherical space which has zero or constant positive curvature, respectively. The negative curvature enables properties such as divergence of parallel lines and exponential volume growth towards the boundary [3]. This volume growth property makes hyperbolic space an ideal candidate for embedding hierarchical and graph structured data [29], and has found many machine learning applications.
Lorentz Model.
The Lorentz model is a way to represent a hyperbolic space. It frames the $d$-dimensional hyperbolic space $\mathbb{H}^{d}$ with curvature $c$ within an $(d+1)$-dimensional Euclidean space $\mathbb{R}^{d+1}$ as follows:
$\mathbb{H}^{d}=\left\{\mathbf{x}\in\mathbb{R}^{d+1}\mid\langle\mathbf{x},\mathbf{x}\rangle_{\mathbb{H}}=-1/c,x_{0}>0\right\}\ ,$ (1)
where the Lorentzian inner product is defined as,
$\langle\mathbf{x},\mathbf{y}\rangle_{\mathbb{H}}=-x_{0}y_{0}+\sum_{i=1}^{d}x_{i}y_{i}\ .$ (2)
Here, the 00-th dimension of the vector is treated as the time component and the rest as the space component. From the definition of $\mathbb{H}^{d}$, the time component can be written using the space component as follows:
$x_{\text{time}}=x_{0}=\sqrt{1/c+\|\mathbf{x}_{\text{space}}\|^{2}}\ ,$ (3)
[…]
Geodesics.
Geodesics in hyperbolic space are the shortest paths between points, analogous to straight lines in Euclidean geometry. In the Lorentz model, geodesics are the intersections of planes through the origin with the hyperboloid.
Tangent Spaces.
The tangent space at a point $\mathbf{x}\in\mathbb{H}^{d}$ in hyperbolic space, is a Euclidean space that locally approximates hyperbolic space around $\mathbf{x}$. Exponential and logarithmic maps are used to project a point from a tangent space to hyperbolic space and vice versa.
Figure 2: An example hierarchy tree. a) Image $I$ with object-level bounding boxes. Each bounding box is entailed by $I$. b) Hierarchy trees created via bounding box-to-bounding box entailment within $I$ (larger bounding boxes entail smaller ones). c) Cross-image hierarchy created by sampling $N$ bounding boxes (with corresponding object classes) from other images, which are then entailed by $I$.
3 Enforcing User-Defined Hierarchies
Our aim is to define a hierarchy in images and enforce it in the latent space using hyperbolic geometry. We first define a part-based hierarchy in images, then discuss our approach to enforce it, and finally introduce our hierarchical retrieval metric.
3.1 Part-Based Image Hierarchy
Visual hierarchies can be established in different ways, depending on the application. In this work, we are interested in a hierarchy that encapsulates the semantic relationship among objects in a scene. For this, scene-object-part hierarchy is appealing as it is useful for applications such as fine-grained object retrieval, object localization, and general scene understanding. This hierarchy has also been shown to emerge in hyperbolic image embeddings [21, 37].
Given an image dataset with bounding box and object class annotations, we define a part-based image hierarchy where the full scene images – typically containing multiple objects – represent the highest level in the hierarchy, while individual objects constitute progressively lower levels, entailed by the full image. In this, larger bounding boxes that significantly envelope smaller ones are considered to entail those smaller ones, establishing a nested hierarchy. In this way, from the full scene to the smallest bounding box, a hierarchy can be defined by recursively applying the entailment rule: if $B$ contained in $A$, then $A$ entails $B$, denoted as $A\to B$. An illustrative example is shown in Fig. 2, where an example hierarchy tree could be road scene $\to$ cyclist $\to$ bicycle $\to$ wheels.
Pairwise Entailment.
Our entailment rule above naturally facilitates a pairwise relationship. Let $I\in\mathcal{I}$ be an image, either the full scene image or a cropped bounding box, and let $\mathcal{B}$ denote the set of all bounding boxes in the dataset. Suppose $\mathcal{B}_{I}$ be the set of bounding boxes contained in the image $I$, i.e, $\mathcal{B}_{I}=\{b\in\mathcal{B}\mid b\subset I\}$ [^1] We use the $\subset$ notation to denote contained in relationship. In the case where $I$ is a cropped bounding box, this relationship is defined to hold if the majority (e.g., 80%) of the small bounding box $b$ is contained in $I$.. Note, each bounding box has an associated object label denoted by $b_{l}\in\mathcal{L}$.
We define the following pairwise entailment relationship:
$\operatorname{if}\ \ b\in\mathcal{B}_{I},\ \ \operatorname{then}\ \ I\to b\ .$ (4)
By applying this recursively, hierarchy trees can be formed as shown in Fig. 2b.
Furthermore, if $I$ is a full scene image, we define an additional entailment relationship across images at the object level. Specifically, for each bounding box $b$ in the image $I$, we sample $K$ bounding boxes from other images with the same label $b_{l}$ and enforce entailment with the image $I$. Formally, let $a\in\mathcal{L}$, and let $\mathcal{B}^{a}_{I,K}\sim\{b\in B\mid b\not\subset I,b_{l}=a\}$ be the set of $K$ bounding boxes of label $a$ on images other than $I$. We enforce the entailment as follows:
$\operatorname{for}\operatorname{all}\ \ b\in\mathcal{B}_{I},\ \ \operatorname{if}\ \ x\in\mathcal{B}^{b_{l}}_{I,K},\ \ \operatorname{then}\ \ I\to x\ .$ (5)
This additional entailment across images reinforces the semantic link between scenes and similar objects across different images (see Fig. 2c).
We posit that these relationships help to structure a hierarchical understanding of images based on scene, object, and part relationships.
Hierarchy Tree.
As noted above, the part-based hierarchy is a tree structure and given an image (or a cropped bounding box), the tree can be traversed using the pairwise entailment relationships. For evaluation, we construct this hierarchy tree per scene/object on the full training set. This hierarchy tree would be used to evaluate the hierarchical image retrieval results. For instance, when considering parent-to-child retrieval, the tree would be traversed from the query image to its leaves and vice versa. More implementation details can be found in supplementary materials.
A subtle point to note is that, in Eq. 4, all the bounding boxes have the full image as a parent node. So if there are other enclosing bounding boxes, then a single bounding box can have multiple parent nodes. In this case, the parent node at the most granular level is chosen and other links are removed when constructing the tree. However, when training our model, the hierarchy tree is not used and only the pairwise entailment relationships are enforced. As we will clarify in the next section, our loss has the flexibility to arrange the embeddings to satisfy this tree structure without violating the pairwise entailment.
3.2 Angle-Based Entailment Loss
We require an asymmetric distance function to enforce pairwise entailment relationships in hyperbolic space. To this end, we adapt the recently proposed hyperbolic-angle-based entailment loss [37]. This is a smooth contrastive variant of the entailment cone loss [11]. It was originally proposed to align text and image modalities with flexibility to “accept” the modality gap along the radial axis. This flexibility is crucial even in our case, where we enforce tree-like hierarchical relationships only via pairwise entailment. The degree of freedom along the radial axis allows our model to structure embeddings in a hierarchical manner without violating the pairwise entailment.
Our loss is a bidirectional version of [37], as illustrated graphically in Fig. 3. In particular, given embeddings of an entailment pair, $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, where $\mathbf{x}$ entails $\mathbf{y}$, we maximize the angles $\beta_{1}$ and $\alpha_{2}$. This enforces entailment in a bidirectional manner. The angles $\beta_{1}$ and $\alpha_{2}$ can be computed using the exterior angle as follows:
$\beta_{1}(\mathbf{x},\mathbf{y})$ $=\pi-\operatorname{ext}(\mathbf{x},\mathbf{y})\ ,$ (6)
$\alpha_{2}(\mathbf{y},\mathbf{x})$ $=\operatorname{ext}(\mathbf{y},\mathbf{x})\ ,$
where $\operatorname{ext}(\mathbf{x},\mathbf{y})$ is the exterior angle between $\mathbf{x}$ and $\mathbf{y}$ and takes the following form [11]:
$\text{ext}(\mathbf{x},\mathbf{y})=\cos^{-1}\left(\frac{y_{\text{time}}+x_{\text{time}}\,c\langle\mathbf{x},\mathbf{y}\rangle_{\mathbb{H}}}{\|\mathbf{x}_{\text{space}}\|\sqrt{\left(c\langle\mathbf{x},\mathbf{y}\rangle_{\mathbb{H}}\right)^{2}-1}}\right)\ .$ (7)
In contrast to [37], in our case an embedding can belong to multiple entailment pairs in a batch. This corresponds to a case of multiple positives in the contrastive loss. Thus, we employ the InfoNCE loss [32] to align all entailment pairs while pushing apart the rest of the negative pairs. Precisely, let $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}$ be a batch of entailment pairs, then the InfoNCE loss for parent-to-child can be written as:
$L^{p\to c}(\mathcal{D},\kappa)=$ (8)
$-\mathbb{E}_{\mathcal{D}}\!\left[\log\frac{\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i})}{\tau}\right)}{\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i})}{\tau}\right)+\sum_{\mathbf{y}^{-}\in\mathcal{N}_{i}}\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i}^{-})}{\tau}\right)}\right]\ ,$
where $\mathcal{N}_{i}$ denotes the set of samples that do not have an entailment relationship with the parent embedding $\mathbf{x}_{i}$. Here, $\kappa:\mathbb{R}^{d\times d}\rightarrow\mathbb{R}$ is the similarity function, and $\tau$ is the temperature parameter. Now, our bidirectional entailment loss can be written as:
$L_{\text{angle}}(\mathcal{D})=L^{p\to c}(\mathcal{D},\beta_{1})+L^{c\to p}(\mathcal{D},\alpha_{2})\ ,$ (9)
Here, the similarity function $\kappa$ is replaced with angles $\beta_{1}$ and $\alpha_{2}$ so that the contrastive loss maximizes angles $\beta_{1}$ and $\alpha_{2}$ for matching entailment pairs in the batch $\mathcal{D}$.
In our implementation, we use a shared image encoder for both parent and child embeddings. Following the reparametrization of [8], we encode the space component of the Lorentz model in the tangent space at origin, use the exponential map to project to the hyperboloid, and the time component is obtained via Eq. 3 to compute the loss.
Loss in Euclidean Space.
This entailment angle loss is general and can be effectively enforced in Euclidean space. In Euclidean space, the exterior angles are formulated as follows:
$\text{ext}(\mathbf{x},\mathbf{y})_{\mathbb{E}}=\cos^{-1}\left(\frac{\lVert\mathbf{y}\rVert^{2}-\lVert\mathbf{x}\rVert^{2}-\lVert\mathbf{x}-\mathbf{y}\rVert^{2}}{2\lVert\mathbf{x}\rVert\cdot\lVert\mathbf{x}-\mathbf{y}\rVert}\right)\ .$ (10)
Now, the loss can be analogously derived.
Figure 3: Learning multi-level hierarchies via contrastive entailment angle loss. Our model first encodes parent-to-child pairs into embeddings, then maximizes $\beta_{1}$ and $\alpha_{2}$ using our contrastive entailment angle loss.
3.3 Hierarchical Retrieval Evaluation
To evaluate retrieval performance on the hierarchy tree, we also introduce a metric that captures the label distribution in the dataset. This is important as different labels can have different numbers of instances and the standard metrics such as Precision@k and Recall@k are agnostic to it.
To this end, consider the parent-to-child relationship, and let $\mathcal{H}_{I}$ denote the hierarchy tree originating from the query image $I$, containing $m$ labels. Then, the labels in $\mathcal{H}_{I}$ are the ground truth labels for the hierarchical retrieval for image $I$. Now, let $\mathbf{h}_{I}\in\mathbf{R}^{m}$ be the precomputed label distribution of $\mathcal{H}_{I}$. Then, our optimal transport (1-D Wasserstein) distance between the retrieved label distribution $\mathbf{r}_{I}$ and $\mathbf{h}_{I}$ is:
$\operatorname{OT}(\mathbf{h}_{I},\mathbf{r}_{I})=\operatorname{Wasserstein}(\bar{\mathbf{h}}_{I},\bar{\mathbf{r}}_{I})\ ,$ (11)
where ¯$\bar{\cdot}\in\mathbf{R}^{m+1}$ is the label distribution with other class added [^2] For $\mathbf{h}_{I}$ other class has zero mass, and for $\mathbf{r}_{I}$ all labels not in $\mathcal{H}_{I}$ are combined to form the other class.. Note that a smaller distance indicates better alignment.
4 Related Work
Hyperbolic geometry allows for exponential volume expansion towards the boundary [3], making it effective for embedding hierarchical structures. For example, hierarchical trees can be embedded with low distortion in hyperbolic space using just two dimensions, which is not possible in Euclidean space even with an unbounded number of dimensions [40]. These advantages have led to significant research into leveraging hyperbolic representations for various data types, including molecular structures [48], 3D [4, 18, 45], images [21, 14, 13, 22, 35], text data [42, 42, 12, 49, 19], and vision-language data [8, 37, 23, 1].
Hyperbolic embeddings can be learned through standard deep learning layers [20] with hyperbolic projection [29] or using hyperbolic neural networks [12]. Many prior NLP, computer vision and knowledge graph studies learn hierarchies from partially order data [12, 43, 27], or minimizing geodesic distance or maximizing similarities [29, 30, 47, 13]. Ganea et al. [12] introduced an angle-based entailment cone loss which pushes child nodes into the cone emanating from the parent node embedding. This approach has been applied to both text [12] and image data [9] with label hierarchies. Recently, this hyperbolic entailment loss was adapted for contrastive learning to develop representations in vision-language models [8, 37, 1, 33]. However, such methods remain relatively unexplored in the image domain. We adapted the angle-based entailment loss from [37] to encode part-based image hierarchies. Unlike previous image models that depend on label hierarchies, such as from WordNet [9], hand labels [28] or simple scene-to-object hierarchies within the same image [13], we capture more complex multi-level part-based hierarchies directly from image data, going beyond visual similarity.
5 Experiments
Datasets.
In order to train and evaluate methods, we construct HierOpenImages, a novel dataset that contains pairwise part-based image hierarchies. It is built from the OpenImages dataset [25] following the steps in Sec. 3.1.
Models.
We evaluate the performance of the proposed method on two popular visual encoders 1) CLIP ViT (B/16) [36], pretrained on large-scale image-text pairs from the Internet, and 2) MoCo-v2 (ResNet-50) [6], pretrained on ImageNet. The embeddings are projected to dimension 128 in the final layer. The hyperbolic model has a learnable curvature parameter. Each model is fine-tuned for a single epoch on HierOpenImages using the contrastive angle-based entailment loss. We denote the model fine-tuned in hyperbolic space as ‘Hyp’, in Euclidean space as ‘Euc’, and the models without fine-tuning as ‘ZeroShot’. Refer to supplementary materials for more experimental details and results.
5.1 Part-based Image Retrieval
Denoting the full image as parent and the bounding box as child, we evaluate retrieval tasks in both child-to-parent and parent-to-child directions. Taking child-to-parent as an example, for fine-tuned models, the retrieval is performed by finding the Top-K parent images corresponding to maximum $\alpha_{2}$ as denoted in Fig. 3. For zero-shot CLIP and MoCo-v2 models, we evaluate results based on maximal cosine similarity, which is consistent with their pre-training. We also evaluated the recent hyperbolic image model, namely HCL [13], on these tasks; however, it did not outperform its backbone zero-shot model, MoCo-v2, in the part-based retrieval task. Please see supplementary materials for more details.
In same-class retrieval, we assess whether the retrieved image belongs to the same class as the query image. In hierarchical retrieval, we verify if the retrieved image exists within the hierarchy tree of the query image, specifically evaluating the quality of the learned hierarchical representations.
Same-Class Retrieval.
In Table 1, we show the top-$K{=}\{5,10,50,100\}$ retrieval accuracy for child-to-parent and parent-to-child. We notice a significant and consistent improvement with the fine-tuned models, in particular the hyperbolic model, across all metrics and model variants. This highlights the relevance of our fine-tuning set-up and the advantages of learning hierarchical part-based image embeddings in hyperbolic space.
Parent-to-child retrieval results show that the performance of fine-tuned models remains stable, while the zero-shot performance declines quickly. This trend likely occurs because zero-shot CLIP and MoCo-v2 rely primarily on visual similarity for retrieval; thus, their performance diminishes when the dataset lacks a sufficient number of visually similar objects for a given query. In contrast, after fine-tuning, the model learns the hierarchical structure of objects, enabling it to retrieve relevant objects even when they have substantially different appearances.
Vision Encoder Model Top-5 Top-10 Top-50 Top-100
Child-to-Parent
Hyp 73.37 72.59 69.84 68.62
Euc 67.83 68.37 67.12 66.04
CLIP ViT ZeroShot 26.73 25.95 23.69 22.70
Hyp 55.53 55.31 51.53 49.51
Euc 54.51 54.16 51.56 49.53
MoCo-v2 ZeroShot 17.23 16.86 16.33 16.07
Parent-to-Child
Hyp 66.02 66.63 66.50 65.91
Euc 65.38 65.70 66.01 65.79
CLIP ViT ZeroShot 47.52 46.60 43.50 42.08
Hyp 52.01 51.64 50.83 50.51
Euc 47.11 46.86 46.95 47.22
MoCo-v2 ZeroShot 18.49 18.37 17.73 17.45
Table 1: Part-based same-class image retrieval evaluation. For child-to-parent image retrieval, the retrieved parent must contain the object class of the query child. For parent-to-child image retrieval, the retrieved child must match a class within the parent.
[…]
CLIP ViT
Hyp 47.28 74.06 16.36
Euc 46.32 72.96 17.00
K=150,000 ZeroShot 36.45 51.42 24.30
Hyp 42.37 82.34 19.27
Euc 41.47 80.98 20.00
K=200,000 ZeroShot 35.61 64.27 29.94
Hyp 38.37 88.79 22.21
Euc 38.09 87.97 22.45
K=250,000 ZeroShot 35.21 76.91 25.29
MoCo-v2
Hyp 46.70 56.09 12.41
Euc 45.70 55.81 13.72
K=150,000 ZeroShot 38.34 48.07 23.41
Hyp 43.81 71.61 15.29
Euc 43.07 71.86 16.83
K=200,000 ZeroShot 37.18 62.56 24.63
Hyp 40.41 83.13 19.09
Euc 39.99 83.42 20.17
K=250,000 ZeroShot 36.24 76.77 25.65
Table 2: Hierarchical evaluation of parent-to-child image retrieval. OT distance is defined in Sec. 3.3.
Hierarchical Retrieval.
Table 2 evaluates the hierarchical structure of the embedding space by retrieving a large number of child images from parent images. We check the alignment between the retrieved distribution and the underlying hierarchical distribution of the test set. Good alignment is a desirable property for fine-grained retrieval as the retrieved set should capture the hierarchies present in the data distribution. The alignment is measured by the optimal transport (Wasserstein distance) between two distributions, with a smaller distance indicating a closer match (see Sec. 3.3). Precision measures the percentage of the retrieved images that are correctly positioned on the lower level of the ground truth hierarchy tree, while recall measures the percentage of ground truth images that are successfully retrieved.
As shown in Table 2, the fine-tuned models in Euclidean and hyperbolic space are significantly improved over the zero-shot CLIP and MoCo-v2 models. The model trained in hyperbolic space better represents the hierarchical distribution of the test set compared to the Euclidean model, outperforming the latter in all cases for precision, and in 4 out of 6 instances for the recall.
Effect of Enforcing Cross Image Entailment.
To evaluate the emergent behaviors of models trained in hyperbolic versus Euclidean spaces, we trained the CLIP ViT model solely on hierarchical part-based entailment data within individual images (entailment pairs with high visual similarity), omitting any cross-image image-to-bounding-box samples.
Different from Top-K retrieval, we compute hierarchical precision and recall metrics as shown in Fig. 5, by applying progressively increasing angle thresholds for $\beta_{1}$ (range 00 to $\pi$) in the fine-tuned models and cosine similarity thresholds (range 0 to 1) for the zero-shot CLIP model to generate the Precision-Recall (PR) curve. As shown, the model fine-tuned in hyperbolic space without cross-image samples outperforms the Euclidean model by a large margin and is even better than the model trained with additional cross-image samples. This clearly indicates the hyperbolic model learns the structure implicitly. When both models are trained with cross-image samples, the hyperbolic model still exhibits better performance than the Euclidean model, however the performance gap is reduced.
Figure 4: Example of parent-to-child retrieval using CLIP ViT zero-shot and fine-tuned model. Results are ordered by embedding norms. Our model retrieves images matching the predefined scene-object-part hierarchy, beyond visual similarity. Best viewed when zoomed in.
Qualitative Results.
In Fig. 4, we show examples of parent-to-child retrieval, with results ordered by increasing embedding norm. The hyperbolic model effectively learns the predefined scene-object-part hierarchy, placing high-level objects near the origin. It captures diverse, visually different objects that might appear in the scene, even if absent from the query image. In contrast, the zero-shot model struggles to link large scenes with smaller objects.
Figure 5: Precision-Recall curves of zero-shot and fine-tuned CLIP ViT models. Dotted lines show models trained solely on hierarchical part-based entailment data within the same images; solid lines represent models trained with additional cross-image scene-to-object samples. Angle or cosine similarity threshold values are marked by text.
5.2 Zero-shot Image Classification
We evaluate the generalization properties of the fine-tuned CLIP model on zero-shot image classification task among 10 popular datasets as shown in Table 3. Consistent with the retrieval results, our model shows significant improvement in zero-shot image classification accuracy on 8 of 10 datasets, with the hyperbolic model consistently outperforming the Euclidean model. This indicates that the learned hyperbolic embeddings effectively capture the underlying structure and relationships within the data, allowing for stable and reliable predictions. This also suggests that the latent space is both well-conditioned and generalizable, as we observe consistent performance across diverse datasets.
Model
CIFAR-10 [24]
CIFAR-100 [24]
SUN397 [46]
Caltech-101 [26]
STL-10 [10]
Tiny-ImageNet [7]
Food101 [2]
CUB [44]
Pets [34]
Flowers [31]
Hyp 84.1 49.7 38.3 79.7 94.7 33.3 42.3 20.5 34.4 36.9
Euc 80.2 44.8 31.8 78.3 93.1 29.2 23.8 4.8 11.6 7.0
ZeroShot 55.1 28.7 39.7 62.2 63.5 11.8 15.0 14.3 31.6 40.0
Table 3: Zero-shot image classification evaluation of CLIP ViT (Top-5 majority vote). Accuracy is in percentage.
6 Conclusion
In this work, we introduce a new learning paradigm that effectively encodes user-defined multi-level visual hierarchies in hyperbolic space without requiring explicit hierarchical labels. We present a concrete example of defining a multi-level part-based image hierarchy using object-level annotations and propose a contrastive loss in hyperbolic space to enforce pairwise entailment relationships. Additionally, we introduce new evaluation metrics for hierarchical image retrieval. Our experiments demonstrate our model effectively learns the predefined image hierarchy and goes beyond visual similarity.
APPENDICES
Appendix A Hierarchy Tree
Our part-based image hierarchy framework is designed to be widely applicable to general image datasets with bounding box annotations. In this section, we outline the key implementation details involved in constructing hierarchy trees.
We introduce a general method to generate dataset-specific ground truth hierarchy trees based on data statistics. Following the approach outlined in Sec. 3.1 of the main paper, we begin by identifying bounding box pairs with substantial overlap. These pairs are then filtered based on two criteria: a frequency threshold and a proportion threshold. For each pair, we record the frequency of occurrences (e.g., bicycle-to-wheel relationships) and calculate the proportion of instances where a child class appears within a parent class (e.g., the percentage of bicycle bounding boxes containing a wheel bounding box). Only pairs meeting both criteria, frequent occurrence and consistent labeling, are preserved. We choose $\textit{frequency}=50$ and $\textit{proportion}=10\%$ in our experiments.
Once entailment pairs are established, they are organized into hierarchical trees (see examples in Fig. 6). In the evaluation of hierarchal image retrieval, the order of the hierarchy tree is essential: for parent-to-child retrieval, lower-level concepts below the child in the hierarchy tree are considered correct, while for child-to-parent retrieval, higher-level concepts above the input classes are correct.
Figure 6: Example of a subset of hierarchical trees extracted from the OpenImages dataset.
Appendix B Experiment Details
B.1 Hyperparameters and training details
We employ the AdamW optimizer with parameters $(\beta_{1},\beta_{2})=(0.9,0.999)$ and a learning rate of $2\times 10^{-5}$. Training was conducted using 8 $\times$ A10G Nvidia GPUs. For each model, we used the largest batch size that fit in memory: CLIP ViT was trained with a total effective batch size of 320, and MoCo-v2 with a total effective batch size of 800. Each model was fine-tuned for a single epoch on HierOpenImages dataset, taking approximately 26 hours for CLIP ViT and 18 hours for MoCo-v2.
During training, we filter out bounding boxes that occupy less than $1\%$ of the full image area, as well as pairs involving small bounding boxes labeled as ‘IsGroupOf’ objects in the bounding box-to-bounding box relationships. For data augmentation, we apply randomly horizontal flip ($20\%$), vertical flip($20\%$), rotate (degree =15), color jitter (brightness=0.2, contrast=0.2, saturation=0.2, hue=0.1), Gaussian blur (kernel size=5, $\sigma=(0.3,1.5)$), and then resize each image to $224\times 224$.
B.2 Part-based Image Retrieval
Data.
HierOpenImages is built from the OpenImages dataset [25], which originally contains approximately 1.9 million images, 14 million bounding boxes and 600 labels. We create image-to-bounding box pairs, including one cross-image bounding box sample for each bounding box class in the image, and bounding box to bounding box pairs where at least $80\%$ of the smaller bounding box is contained in the larger bounding box.
The HierOpenImages test set contains approximately 1 million full images and 3 million bounding box images. During part-based image retrieval evaluation, to limit the number of object classes within a bounding box image, we only evaluate on bounding box images that occupy 5%-30% of the full image area. This reduced the candidate full images and bounding box images to 0.33M and 0.6M, respectively.
For part-based image retrieval, we randomly select a subset of 10,000 full images and 10,000 bounding box images as query images, using the entire test set as candidates. The bounding box images in the query set span 339 object classes, while the full images are labeled with 516 classes. The top-50 class frequency distributions for both the query and candidate sets are shown in Fig. 7. Although the query and candidate set distributions are similar, the class distribution is highly imbalanced, highlighting the importance of hierarchical retrieval evaluation using combined precision-recall curves and OT distances (Sec. 3.3; see Fig. 5 and Table 2 in the main paper).
Figure 7: Bounding box class distribution in the candidate and query sets.
Same-Class Retrieval.
For part-to-full retrieval, a retrieval is considered correct if the retrieved full image contains the same object class as the query bounding box image. For full-to-part retrieval, a retrieval is correct if the retrieved bounding box image corresponds to an object class within the query full image.
Metric Model Top-5 Top-10 Top-50 Top-100
Child-to-Parent
Cos Sim. HCL 15.29 15.34 15.12 14.95
Hyp Dist. HCL 14.64 14.72 14.45 14.27
Hyp Angle Ours 55.53 55.31 51.53 49.51
Parent-to-Child
Cos Sim. HCL 17.10 16.90 16.45 16.23
Hyp Dist. HCL 16.72 16.52 16.20 16.00
Hyp Angle Ours 52.01 51.64 50.83 50.51
Table 4: Part-based same-class image retrieval evaluation on HCL [13] model and our model.
[…]
Cos Sim.
K=150,000 HCL 36.05 46.29 24.46
K=200,000 HCL 35.82 61.38 24.87
K=250,000 HCL 35.62 76.43 25.24
Hyp Dist.
K=150,000 HCL 35.60 45.39 25.10
K=200,000 HCL 35.32 60.22 25.63
K=250,000 HCL 35.15 75.23 25.96
Hyp Angle
K=150,000 Ours 46.70 56.09 12.41
K=200,000 Ours 43.81 71.61 15.29
K=250,000 Ours 40.41 83.13 19.09
Table 5: Hierarchical evaluation of parent-to-child image retrieval on HCL [13] model and our model. The OT distance is defined in Sec. 3.3 in the main paper.
Model
CIFAR-10 [24]
CIFAR-100 [24]
SUN397 [46]
Caltech-101 [26]
STL-10 [10]
Tiny-ImageNet [7]
Food101 [2]
CUB [44]
Pets [34]
Flowers [31]
Query N 50,000 50,000 1,829 1,829 8,000 10,000 25,250 5,794 3,669 6,149
Candidate N 10,000 10,000 7,315 7,315 5,000 100,000 75,750 5,994 3,680 1,020
Table 6: Query and candidate set sizes for each dataset used in zero-shot image classification evaluation.
Hierarchical Retrieval.
From a query parent image, correct child classes are all classes located at the lower level on the hierarchy tree of the labeled classes of the parent image (see examples in Fig. 6). For instance, when querying with a high-level full image, such as an image of a car, we expect to retrieve lower-level bounding boxes associated with the car, such as the car mirror, wheel or car plate etc. Conversely, when querying with a bounding box image, such as a wheel, we expect to retrieve various types of higher-level full images that include wheels, like cars, bicycles or cyclists etc.
Hierarchal Retrieval Evaluation
To compute the optimal transport (1-D Wasserstein) distance between the retrieved label distribution and the ground truth (Table 2 and Fig. 5 in the main paper), we construct the ground truth distribution based on the frequency of each class (and its hierarchical classes) in the query parent image. We count the occurrences of each class in the candidate set and build the ground truth distribution by normalizing the frequencies to sum to 1. Similarly, the retrieval distribution is built by counting and normalizing retrieved class occurrences, and assigning any retrieved classes outside the ground truth hierarchy tree to an ‘others’ class. The two distributions are aligned by class order (ground truth distribution is zero in the ‘others’ class), and the 1-D Wasserstein distance is computed using the scipy library.
Note that the Wasserstein distance has a closed-form formula for 1-D data. If $P$ and $Q$ are represented as discrete empirical distributions (e.g., histograms or sorted samples of size $n$), let $\{x_{1},x_{2},\ldots,x_{n}\}$ to be sorted values $P$, and $\{y_{1},y_{2},\ldots,y_{n}\}$ to be sorted values from $Q$, then the 1-D Wasserstein distance is:
$W_{p}(P,Q)=\left(\frac{1}{n}\sum_{i=1}^{n}\left\|x_{i}-y_{i}\right\|^{p}\,\right)^{1/p},$
where $p$ refers to the order of the distance in the general p-Wasserstein metric.
In Table 2 of the main paper, we retrieve the Top-K results starting from approximately $50\%$ of the test data size. This is necessary because each image contains an average of 5.29 distinct classes and 61.5 classes across hierarchical trees, requiring a large number of retrievals to accurately evaluate the distribution.
B.3 Zero-shot Image Classification
We use the official splits from torchvision, designating the training set as candidates and the test/validation set as queries. For datasets without predefined splits on torchvision, such as SUN397 [46] and Caltech101 [10], we randomly split the dataset into $80\%$ candidates and $20\%$ queries. The sizes of query and candidate images for each dataset are listed in Table. 6.
For each dataset, we retrieve the top-5 most similar images from the candidate set based on cosine similarity and predict the class by majority vote among top images. We chose cosine similarity instead of hyperbolic angles as our metric, as there is no part-based relationship between the query and candidate images, and this is not an evaluation of preserving a pre-defined hierarchy.
B.4 Comparison to recent hyperbolic image model.
We compared our method with the recent hyperbolic image model, HCL [13], on the part-based same-class and hierarchical image retrieval tasks (see Table 4-5). HCL is trained on the OpenImages dataset using scene-to-object relationships and demonstrates hierarchical structure in its embeddings, making it, to our knowledge, the most similar method to ours.
HCL learns scene-to-object relationships by minimizing the hyperbolic distance between scene images and their corresponding object bounding box images, while also maximizing the cosine similarity between slightly shifted and jittered crops of the same object. We evaluated its retrieval performance using both cosine similarity and hyperbolic distance metrics. We compare our MoCo-v2 hyperbolic model with HCL, as both use the same visual backbone. Instead of fine-tuning, HCL retrained MoCo-v2 on the OpenImages dataset. HCL underperformed compared to the pretrained MoCo-v2 model trained purely in Euclidean space (see Table 1-2 in the main paper), and performed significantly worse than our model on these tasks. These results indicate that while HCL is conceptually similar, it is not a suitable comparison algorithm for our task.
Appendix C Ablation Studies
In this ablation study, we evaluate the impact of cross-image scene-to-object samples. We fine-tuned the CLIP ViT model solely on hierarchical part-based entailment data within individual images (entailment pairs with high visual similarity), excluding any cross-image image-to-bounding-box samples. Fig. 5 in the main paper shows that the hyperbolic model fine-tuned without cross-image samples significantly outperforms the Euclidean model and even surpasses the model trained with additional cross-image samples. This indicates that training in hyperbolic space enhances the model’s ability to recognize visually dissimilar entailment pairs.
In this supplementary material, we further evaluate these models on part-based same-class and hierarchical image retrieval tasks, as shown in Table. 7-8. The best results are highlighted in bold and the second-best results are shown in blue. The results clearly indicate that cross-image sampling improves image retrieval performance. Notably, the results align with Fig. 5 in the main paper, with the hyperbolic model trained without cross-image samples outperforming the Euclidean model trained with cross-image samples in 11 out of 17 cases.
Model Cross-Image Top-5 Top-10 Top-50 Top-100
Child-to-Parent
Hyp 73.37 72.59 69.84 68.62
Euc 67.83 68.37 67.12 66.04
Hyp 70.47 70.29 67.89 66.38
Euc 64.98 64.49 63.51 63.10
ZeroShot - 26.73 25.95 23.69 22.70
Parent-to-Child
Hyp 66.02 66.63 66.50 65.91
Euc 65.38 65.70 66.01 65.79
Hyp 63.00 63.78 63.65 63.44
Euc 60.33 60.73 61.43 61.38
ZeroShot - 47.52 46.60 43.50 42.08
Table 7: Part-based same-class image retrieval evaluation on fined-tuned and zero-shot CLIP ViT models. Cross-image ✓indicates models fine-tuned on entailment relationships both within and across images at the category level, while ✗ represents models fine-tuned without cross-image sampling. The evaluation setup is the same as Table 1 in the main paper. The best results are marked in bold, and the second-best results are in blue.
[…]
Hyp 47.28 74.06 16.36
Euc 46.32 72.96 17.00
Hyp 46.87 73.52 16.95
Euc 44.93 71.34 18.43
K=150,000 ZeroShot - 36.45 51.42 24.30
Hyp 42.37 82.34 19.27
Euc 41.47 80.98 20.00
Hyp 41.94 81.58 19.77
Euc 40.02 78.89 21.25
K=200,000 ZeroShot - 35.61 64.27 29.94
Hyp 38.37 88.79 22.21
Euc 38.09 87.97 22.45
Hyp 38.00 88.10 22.55
Euc 36.87 86.04 23.53
K=250,000 ZeroShot - 35.21 76.91 25.29
Table 8: Hierarchical evaluation of parent-to-child image retrieval on fined-tuned and zero-shot CLIP ViT models. The OT distance is defined in Sec. 3.3 (main paper), and the evaluation setup follows Table 2 (main paper). The best results are marked in bold, and the second-best results are in blue.
Appendix D More Qualitative Results
More qualitative parent-to-child retrieval results are shown in 8. Bounding box images are filtered by angle (hyperbolic model) or cosine similarity (zero-shot model) thresholds and sorted by increasing embedding norms. Our hyperbolic model retrieves diverse and visually distinct lower-level objects related to the query images, organized according to the predefined scene-object-part hierarchy in the embedding space.
Figure 8: Example of parent-to-child retrieval using CLIP ViT zero-shot and fine-tuned model. Results are ordered by embedding norms. Our model retrieves images matching the predefined scene-object-part hierarchy, going beyond visual similarity. All retrieved bounding box images are scaled to the same ratio. Best viewed when zoomed in.